OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = −32 and a9 = −120?

OpenStudy (anonymous):

Do you know how to get the formula?

OpenStudy (anonymous):

I am lost I don't remember

OpenStudy (anonymous):

an = a1 + (n-1)d

OpenStudy (anonymous):

a1 is the first term, d is the difference, and n is the nth term.

OpenStudy (anonymous):

-120=-32+(n-1) d

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

wat do I do after that

OpenStudy (anonymous):

You would put the first term in as "n"

OpenStudy (anonymous):

So it would actually be a1+(-32-1)d=-120

OpenStudy (anonymous):

Can you do the math, please?

OpenStudy (anonymous):

but a1 =-32 so what do u mean

OpenStudy (anonymous):

We're trying to figure out d first

OpenStudy (anonymous):

a1-33d= a9 a1-a9=33d That's when you plug in the numbers.

OpenStudy (anonymous):

(-32)-(-120)=33d -32+120=33d 88=33d 8/3 = d

OpenStudy (anonymous):

Now we have d. So we use the formula: an=a1(n-1)d

OpenStudy (anonymous):

a32=(-32)(32-1)8/3 Can you do the math?

OpenStudy (anonymous):

Let's see if we get the same answer.